# AI-Powered Urban Transportation Digital Twin: Methods and Applications

### v1

@@ 3.4 Use cases
The value of a DT for urban transportation management ranges from traffic state prediction, spatiotemporal traffic flow forecasting, to urban planning and policy making. Here we focus on two that are particularly important in urban settings.
3.4.1 VRU safety warning: Safety-critical application
@@
Safety-critical applications, making automated life-and-death decisions such as collision avoidance warning between automobiles and pedestrians, need to activate at the precise time and the right moment with bounded latency. Safety-critical systems are often time-critical and reactive, because they need to react to external signals in precise time [69], and require tasks to be executed in a timely manner. Theses systems hold the substantial potential to enhance road safety and save lives. They are, however, risky to test and run, because rare events like automobile collisions are challenging and unethical to replicate. Thanks to emerging technologies in ubiquitous sensing, low-latency high-bandwidth communication, high-speed computing and AI, safety critical applications could potentially be modeled, simulated, processed, and tested in a DT. The applications of DTs on safety-critical scenarios, however, remains understudied, because it necessities short runtime and real-time reaction, posing high requirements for communication and computing technologies, pipeline architecture design, and testing.
We summarize the safety-critical applications to address conflict risks between vehicles and VRUs in Tab. 4, and offer an outlook of safety guarantee related methods in Sec. 5.1.4.
Table 4: Literature review on safety warning to VRUs.
@@
Develops a video-based vehicular BSMs method with lower error and latency that outperforms the cellular vehicle-to-everything (C-V2X) method. N/A 1. End-to-end delay: <100absent100<100< 100  ms
@@ 2. Localization/Speed accuracy
Proposes a system using commercial devices and standard messages for road user communication. The distance between VRUs and nearby entities Latency from VRU to CAM <50absent50<50< 50 ms
@@
[…]
@@ 2. MAC channel utilization
Develops a system with sensing and communication, along with fusion and collision detection algorithms, to predict potential collisions and warn VRUs. The distance between VRUs and nearby entities 1. End-to-end latency <300absent300<300< 300 ms
@@ 2. Granular latency per frame
3.4.2 Networkwide adaptive traffic signal control
Learning methods for network wide adaptive traffic signal control (ATSC), including reinforcement learning (RL) and federated learning (FL), aim to learn optimal traffic signal control strategies in traffic environments with unknown state transitions and unknown rewards. Assuming that traffic signal controllers are intelligent autonomous agents, their goal is to make sequential decisions to optimize an objective function, such as minimization of traffic delay or maximization of traffic throughput. To train an optimal policy for traffic signal controllers, a traffic simulator or DT, needs to be built, where dynamic traffic is simulated in the training environment aligned with the testing environment in principle. The trained policies leveraging such a DT need to be executed in the real world, which triggers updated environments for future decision making. Related work are summarized in Tab. 5.
Table 5: Summary of papers using (F)RL for ATSC
Ref. RL FL DT Simulator Real-world Demand Real-world Network
[63] DDPG SUMO
[30] SUMO
[71] A2C Vissim
[64] Vissim
[201] SUMO
[42] A3C FedAvg SUMO
[182] DQN FedAvg Cityflow
[58] PPO FedAvg SUMO
[16] DQN FedAvg SUMO
[109] A2C SUMO
[77] A2C SUMO
[27] A2C SUMO
[52] PPO SUMO
FedAvg: Federated Averaging, DDPG: Deep Deterministic Policy Gradient, A2C: Advantage Actor Critic, A3C: Asynchronous Advantage Actor Critic, PPO: Proximal Policy Optimization, DQN: Deep Q Learning.
3.5 Close the loop
To realize the value of a DT, the feedback from the digital to the physical is the key. The type of feedback and to what entities would define the technological constraints.
  • Feedback to VRUs (super local): When users react to an issued warning, the state of the environment for the VRU would change, and the DT could update its status about whether continuous warning is needed. Normally the warning is safety related, thus demanding low latency and high bandwidth.
  • Feedback to cars (super local): The feedback to cars includes hierarchies: (1) strategic driving feedback, such as routing recommendations, allowing higher latency and lower bandwidth; (2) tactical feedback on tactical driving decisions, such as lane-change, with medium latency and bandwidth; and (3) operational feedback such as throttle control, demanding (ultra) low latency and bandwidth. When cars change their driving maneuvers, the traffic environment needs to be updated and unsafe outcomes could be potentially avoided.
  • Feedback to traffic operators and infrastructure (local or networkwide): For traffic operators, DTs could provide anomaly detection for diagnosis, and predict future traffic states or collisions for prognosis, which would help with variable speed limit, dynamic speed advisory, lane usage adaptation, ramp metering, traffic signal timing optimization, and network wide traffic management, with increasing requirements for latency and bandwidth.
  • Feedback to city planners (citywide): In the long-term, a DT could provide aggregate traffic information to city planners to assess the potential societal impact of new infrastructure planning. Thus, the technological requirement is relatively low, but the collected data needs to cover a relatively large spatiotemporal range for citywide decision making,
3.6 Communication and networking
3.6.1 Real-time requirements and low latency targets
Smart city applications can be grouped according to their latency requirements. Many, if not all, pedestrian-associated application (facilitated by pedestrian detection/observations and message notifications) are likely to expect the round trip delay in the range of a couple of seconds. Such latency can be supported by contemporary cameras, communication protocols and inferencing engines. Applications which would attempt to close the observation/notification loop for vehicles moving at about 10kmph may expect latency in tens of millisecond. Using conventional video compression, RTP/RTSP streaming and edge computing is inadequate to support such latency. This presents the opportunity to pursue novel engineering solutions and research problems.
3.6.2 Communication techniques and protocols
Message Queuing Telemetry Transport (MQTT) is a lightweight communication protocol that offers efficient, reliable, and scalable data transfer necessary for real-time, safety-critical applications. Such a protocol can efficiently collects and distributes traffic data from numerous sensors to update DTs in real-time, and facilitates fast communication of incidents and emergencies, allowing timely interventions. Key features include [122]:
  • Lightweight and scalable: MQTT is designed with a minimal fixed header of only 2 bytes, which reduces the communication overhead significantly. It decouples message producers and consumers, allowing easy scalability and flexibility in handling numerous sensors and devices.
  • QoS levels and reliability: MQTT’s different QoS levels enhance reliability by ensuring message delivery according to the application’s criticality. QoS 0 delivers messages at most once without acknowledgment. QoS 1 ensures messages are delivered at least once, while QoS 2 guarantees exactly one delivery through a four-step handshake, preventing duplication and ensuring data integrity for critical communications. Moreover, MQTT’s LWT (Last Will and Testament) feature allows the system to detect and respond to unexpected device disconnections, maintaining the robustness of the network.
4 Proposed DT pipeline
Figure 2: Architecture of the proposed DT pipeline: Urban Hybrid Twin.
In this section, we present a DT architecture called “Urban Hybrid Twin,” based on the sensing/communication/computing testbed deployed in NYC (Fig. 1). The proposed architecture of “Urban Hybrid Twin” is enabled by CPS, user software defined radios (SDR), cameras and LiDARs, high speed communications, and AI computing deployed in the edge cloud.
Figure 3: Diagram of intersection safety warning use case.
4.1 System components
4.1.1 Physical Infrastructure
The physical functionality of the proposed DT is built on top of the COSMOS testbed (“Cloud enhanced Open Software defined mobile wireless testbed for city-Scale deployment” [125] developed for real-world research, development, and deployment of city-scale advanced wireless communications and innovative applications [125]. It targets the technology “sweet spot” of ultra-high bandwidth and ultra-low latency, a capability that will enable a broad class of safety-time critical applications. Deployed in West Harlem, NYC, next to Columbia University campus, the COSMOS testbed is an enabler for research on the design, development, and deployment of a DT for urban traffic management.
Pilot DT experiments are executed at the intersection of Amsterdam Avenue and 120th Street, where COSMOS cameras, LiDAR, and wireless sensing and communication nodes are deployed. The cameras employ H264 encoding with an I-frame interval set to 10 frames. The live RTSP video streams–with 4K resolution at 30 fps–were processed on a COSMOS edge server equipped with an A100 GPU. The LiDAR is a Velodyne VLP-32C with up to 200m range.
Note that there is a growing trend of studies relying on emerging communication protocols like C-V2X for vehicle DTs [145, 157], which requires customized infrastructure and user-end devices. Our initial experiments utilize contemporary communications and sensing protocols for rapid deployment. They will be replaced by emerging methods that can achieve lower latency and higher bandwidths. Software defined features of the COSMOS testbed and the flexibility of radios make it possible to deploy emerging technologies such as C-V2X and experiment with novel low-latency applications.
4.1.2 Digital twin
4.2 System functionality
4.2.1 Resource optimization
@@
4.2.2 Use cases
4.2.2.1   Pedestrian-vehicle safety warning
Leveraging existing sensors at the intersection, we have developed a combined VRU and vehicle warning application. The cameras and the LiDAR sense the presence of VRUs at urban intersections, make predictions of their movements, apply traffic operation and control strategies, and feedback to system controllers and road users, with the primary goal of increasing traffic safety and efficiency. (see Fig. 3).
A variety of AI models have been developed to predict future trajectories of interacting road users at the intersection, including InfoSTGCAN (An Information-Maximizing Spatial-Temporal Graph Convolutional Attention Network) integrating InfoGAN and spatiotemporal graph convolutional attention network [131], and PI-NeuGODE using physics-informed graph neural ordinary differential equations[107, 110], and uncertainty quantification (UQ) to characterize the predictive confidence [105].
4.2.2.2   Learn to optimize adaptive traffic signal control
To optimize traffic signal control over a road network requires first to predict traffic patterns on road segments and intersections. The AI models we developed include PINO (physics-informed neural operator) [45] that predicts future traffic dynamics by training a mapping from initial conditions to traffic states in time and space, building on our previous work on PIDL [138, 139] and the UQ problem [108, 106]. For ATSC, we have developed an RL method, CVLight [109], and an FRL method [42] on the COSMOS network [42]. Their performances are illustrated in Fig. 4.
Figure 4: (a): ATSC for a traffic corridor. (b): Performance metrics of ATSC methods. Left: travel time and communication cost; Right: traffic volume on the incoming lane over time.
4.3 Evaluation
@@
4.3.1 End-to-end latency
Measuring latency could be particularly problematic due to the need for cross-network timing synchronization of devices and compute server. Spot measurements of several components are shown in Table 6. This table presents the latency incurred by the main components of a typical video analytics pipeline for object detection and tracking, as depicted in Fig. 3, which has been systematically optimized in terms of memory and resource usage. These elements usually run on an edge device or server. The results in Table 6 identify potential bottlenecks within the pipeline and indicate which components could benefit from further optimization. Additionally, the table illustrates the impact of video content complexity on the latency of each component, providing insights into performance variations during busy and sparse traffic conditions.
To understand the sensitivity and complexities of real-time video analytics, we measured the latency introduced by key components of a typical video analytics pipeline, illustrated in Fig. 3. For object detection, we used the YOLOv8m model [149] (batch size: 1) with the Nvidia DCF-based tracker. The YOLOv8m model was configured to accept input images at 832×832 resolution, corresponding to the size of the regions cropped by the RoI cropping component. The 5G and LTE networking latency was measured using iPhone 13 Pro Max, while wired and wireless networks provided by Columbia University Information Technology were used for the Ethernet and WiFi measurements. Latency measurements are conducted at various times of the day with different traffic densities. Tab. 6 summarizes the results of these measurements by comparing the average and standard deviation of latency per frame for the main components of the pipeline.
@@ 4.3.1 End-to-end latency
Reception Pre-processing Object Detection Object Tracking MQTT Msg Creator MQTT Msg Retrieval
Ethernet Wi-Fi LTE 5G
Avg Latency (Std Dev) [ms] 1.94 (1.69) 0.108 (0.024) 7.216 (0.121) 0.973 (0.173) 0.081 (0.021) 3.21 (0.315) 6.86 (1.19) 45.72 (15.30) 39.21 (7.12)
Table 6: Average Latency and Standard Deviation for different pipeline elements.
These latency measurements demonstrate how each pipeline component influences overall latency and its variability under different video content densities. The results highlight the necessity of optimizing these components for real-time applications, particularly in environments with dynamic traffic patterns. In the following, we describe the function of each element in the inference pipeline (see Fig. 3) and their associated latencies based on the results presented in Tab. 6.
  • Reception: Receives RTP (Real-time Transport Protocol) data packets and, after parsing and decoding, converts them to raw video frames. The latency increases under busy traffic conditions, primarily due to higher complexity of the data being processed, such as larger frame sizes and more variation in the number of data packets per frame.
  • Pre-processing: Extracts regions of interest (RoIs) from raw video frames and discards irrelevant regions. It incurs low latency with minimal variations, indicating the effectiveness of pre-processing in optimizing video analytics performance.
  • Object detection: Applies the object detection model to the cropped RoIs. This element incurs the highest latency in the pipeline. The latency is slightly higher in busy traffic/crowd conditions due to more objects in each video frame.
  • Object tracking: Assigns unique IDs to detected objects across video frames for tracking. It has higher latency under busy conditions due to the increased number of objects being tracked simultaneously.
  • MQTT message creator: Converts the extracted metadata into MQTT format [104] and transmits it over a designated topic. Its latency is consistently low, with a slight increase during busy conditions due to the need to assemble larger messages.
  • MQTT message retriever: A web application captures round-trip time (RTT) of MQTT messages, and measures the latency as half an RTT.
4.3.2 Safety warning performance
@@
To validate the accuracy of our trajectory prediction and risk assessment algorithms for warning generation, we conducted three rounds of simulation, each lasting $1010$ minutes and generating a total of $232232$ virtual pedestrians in CARLA. We then compared the number of collision warning messages with the actual number of simulated collisions. The process is illustrated in Fig. 5.
Metrics for risk assessment include time to collision (TTC) (i.e., the time remaining before a collision occurs) and post-encroachment time (PET) (i.e., the time interval between when the encroaching vehicle leaves the conflict point and when the vehicle of the right-of-way arrives at the conflict point). For prediction, we select TTC and set its threshold as $1.21.2$ seconds to minimize both false positives (FP) and false negatives (FN) while optimizing true positives (TP) and true negatives (TN). The resultant confusion matrix is $\text{TP},\text{FP}],[\text{FN},\text{TN}]]=[[66,45],[2,119]].[ [ TP , FP ] , [ FN , TN ] ] = [ [ 66 , 45 ] , [ 2 , 119$ ] ] .
5 Conclusions and open questions
Transportation applications in urban settings are generally challenging to design, develop, deploy, and test for their potential unsafe and unethical consequences. Thus, AI-empowered DT plays a critical role in effective implementation of these applications. Despite relatively sparse literature in this domain, we review an ensemble body of literature on how to leverage emerging technologies in sensing, communication, edge and cloud computing, for urban traffic management. We hope this paper can serve as a pointer to help researchers and practitioners identify challenges and opportunities on the development of DTs; a bridge to initiate conversations across interdisciplinary researchers in sensing, communication and networking, traffic engineering, and AI; a road map to exploiting potentials of DTs for diverse transportation applications.
5.1 Emerging trends

### v2

@@ IV-C Use cases
The value of a DT for urban transportation management ranges from traffic state prediction, spatiotemporal traffic flow forecasting, to urban planning and policy making. Here we focus on two that are particularly important in urban settings, leveraging video analytics.
IV-C1 Use case 1: Micromobility detection
In the U.S., the e-bike had a market size of almost \$2 billion in 2022, with a projected growth of 15.6% from 2023 to 2030. In 2022, 1.1 million e-bikes were sold, four times as many as were sold in 2019. In 2019, 136 million trips were made using micromobility, a 60% increase from 2018. Because their presence and surge have profound implications for road safety and urban traffic management, automatic detection of micromobility for regulation becomes increasingly important. However, there does not exist a standard AI model nor a dataset for the object detection of this emerging transportation mode.
AI model evaluation
[…]
TABLE VI: Class-wise recall (%) on the Micromobility test set. Higher values indicate better detection performance.
Model Bicycle $\uparrow$ e-Bike $\uparrow$ Motorcycle $\uparrow$
YOLOv9e 52.9 54.7
Ours 68.1 61.0 91.8
Figure 6: Four sample frames showing the small scale of micromobility instances from infrastructure-based 4K resolution cameras in the COSMOS testbed.
IV-C2 Use case 2: VRU safety warning
@@
Safety-critical applications, making automated life-and-death decisions such as collision avoidance warning between automobiles and pedestrians, need to activate at the precise time and the right moment with bounded latency. Safety-critical systems are often time-critical and reactive, because they need to react to external signals in precise time [127], and require tasks to be executed in a timely manner. Theses systems hold the substantial potential to enhance road safety and save lives. They are, however, risky to test and run, because rare events like automobile collisions are challenging and unethical to replicate. Thanks to emerging technologies in ubiquitous sensing, low-latency high-bandwidth communication, high-speed computing and AI, safety critical applications could potentially be modeled, simulated, processed, and tested in a DT. The applications of DTs on safety-critical scenarios, however, remains understudied, because it necessities short runtime and real-time reaction, posing high requirements for communication and computing technologies, pipeline architecture design, and testing. We summarize the safety-critical applications to address conflict risks between vehicles and VRUs in Tab. VII, and offer an outlook of safety guarantee related methods in Sec. V-A3.
TABLE VII: Literature review on safety warning to VRUs.
@@
Develops a video-based vehicular BSMs method with lower error and latency that outperforms the cellular vehicle-to-everything (C-V2X) method. N/A 1. End-to-end delay: <100<100  ms
@@ 2. Localization/Speed accuracy
Proposes a system using commercial devices and standard messages for road user communication. The distance between VRUs and nearby entities Latency from VRU to CAM <50<50 ms
@@
Develops a system to deliver CAM to VRUs on smartphones using Beacon stuffing without the need for root access to utilize 802.11p. N/A 1. End-to-end latency $\sim$ 2500 ms
@@ 2. MAC channel utilization
Develops a system with sensing and communication, along with fusion and collision detection algorithms, to predict potential collisions and warn VRUs. The distance between VRUs and nearby entities 1. End-to-end latency <300<300 ms
@@ 2. Granular latency per frame
Intersections, where sixty percent of crashes happen [135, 136], are critical bottlenecks of an urban transportation network. To improve urban road safety and increase traffic capacities, safety warning is the key. Leveraging existing sensors at the intersection, we have developed a combined VRU and vehicle warning application. The cameras and the LiDAR sense the presence of VRUs at urban intersections, make predictions of their movements, apply traffic operation and control strategies, and feedback to system controllers and road users, with the primary goal of increasing traffic safety and efficiency (see Fig. 3).
Resource optimization
@@
In addition to dynamic adaptation of the system’s configuration, we can reduce latency and GPU consumption by identifying Regions of Interest (RoIs) where pedestrians might be in danger, using lightweight processing (e.g., low resolution, small models). As shown in Table VIII, smaller models (“YOLOv8s” in Column 4) incur significantly lower latency. Larger models (“YOLOv8x” in Column 6) with higher resolution and frame rate are only triggered when a critical danger area, i.e., RoI, is detected. For example, detecting large objects such as vehicles does not require large models or high resolution. Therefore, we can use a smaller model and lower resolution to detect vehicles and their trajectories, and based on that, determine the danger areas where pedestrians may be at risk or in the blind zones of vehicles. These identified danger areas are then processed using larger models and higher resolution to detect pedestrians at risk and notify them or the vehicles if necessary. In our intersection safety warning system (shown in Fig. 3), we have embedded an RoI Specifier element to facilitate this approach.
AI models
A variety of AI models are developed to predict future trajectories of interacting road users at the intersection, including InfoSTGCAN (An Information-Maximizing Spatial-Temporal Graph Convolutional Attention Network) [137], which encodes road user trajectories into quantized latent codes to account for heterogeneity in road users, PI-NeuGODE (Physics-Informed Graph Neural Ordinary Differential Equations), using physics-informed graph neural ordinary differential equations [138, 139], and uncertainty quantification (UQ) to characterize the predictive confidence [140]. These probabilistic trajectory prediction methods enhance pedestrian safety at intersections by incorporating UQ into risk assessment. The Kalman Filter (KF) [141] performs prediction by recursively estimating the state of a moving object using a linear dynamic model with Gaussian noise. Trajectron++ [142] models multimodal future trajectories by combining recurrent neural networks with conditional variational autoencoders, enabling probabilistic and socially-aware predictions. We present the predicted trajectories and performance comparison in Fig. 7, where the performance is evaluated using Average Displacement Error (ADE, i.e., the mean Euclidean distance between predicted and ground-truth positions over all future time steps) and Final Displacement Error (FDE, i.e., the Euclidean distance at the final prediction step).
Method ADE $\downarrow$ FDE $\downarrow$
Kalman Filter 0.91 1.92
Trajectron++ 0.77 1.34
ADE and FDE comparison. Lower values are better ($\downarrow$).
Figure 7: Performance comparison for trajectory prediction
Evaluation
@@
a. End-to-end latency
Measuring latency could be particularly challenging due to the need for cross-network timing synchronization of devices and compute server. Table VIII presents the latency incurred by the main components of a typical video analytics pipeline for object detection and tracking, as shown in Fig. 3. We measured the latency for three sizes of YOLOv8 object detection model: YOLOv8s (small, $\sim$11.2 million parameters), YOLOv8m (medium, $\sim$25.9 million parameters), and YOLOv8x (large, $\sim$68.2 million parameters). Larger models offer improved detection accuracy but come at the cost of increased latency and higher GPU and memory usage [143]. The pipeline has been systematically optimized in terms of memory and resource usage. These elements usually run on an edge device or server. The results identify potential bottlenecks within the pipeline and indicate which components could benefit from further optimization.
TABLE VIII: Average latency and Standard Deviation for different pipeline elements.
@@
Reception Pre-processing Object Detection Model Object Tracking MQTT Msg Creator MQTT Msg Retrieval
Small (YOLOv8s) Medium (YOLOv8m) Large (YOLOv8x) Ethernet Wi-Fi LTE 5G
Avg Latency (Std Dev) [ms] 1.94 (1.69) 0.108 (0.024) 4.034 (0.084) 7.216 (0.086) 11.140 (1.800) 0.973 (0.173) 0.081 (0.021) 3.21 (0.315) 6.86 (1.19) 45.72 (15.30) 39.21 (7.12)
b. Safety message generation accuracy
@@
To validate the accuracy of our trajectory prediction and risk assessment algorithms for warning generation, we conducted three rounds of simulation, each lasting $10$ minutes and generating a total of $232$ virtual pedestrians in CARLA. We then compared the number of collision warning messages with the actual number of simulated collisions. The process is illustrated in Fig. 8.
(a) TTC (in seconds)
(b) Distance (in pixels)
Figure 9: ROC curves for trajectory prediction to determine optimal thresholds
Note that metrics for risk assessment include time to collision (TTC) (i.e., the time remaining before a collision occurs) and post-encroachment time (PET) (i.e., the time interval between when the encroaching vehicle leaves the conflict point and when the vehicle of the right-of-way arrives at the conflict point). TTC is commonly used, while PET is typically for post-event analysis. To select the optimal thresholds of TTC, we first generate the ROC (Receiver Operating Characteristic) curve in Fig. 9. Fig. 9(a) illustrates the relationship between the true positive rate (TPR) and the false positive rate (FPR) across various TTC threshold values, ranging from 0.1 to 1.2 seconds. The optimal TTC threshold is 1.1 seconds (indicated by a blue square), at which the TPR reaches $0.958$ and the FPR is $0.4$. To compute TTC, we need to compare each pair of predicted trajectory points in the next $t$ time steps, determined by a danger distance threshold that is defined as when the proximity between a vehicle and a pedestrian constitutes a dangerous interaction. Fig. 9(b) presents the TPR and FPR for the threshold of danger distances ranging from $5$ to $100$ pixels. The optimal threshold is valued at $30$ pixels, which yields a TPR of $0.958$ and an FPR of $0.345$. Based on the above two thresholds, we run the collision prediction model. The resultant confusion matrix for predicted collisions under the selected thresholds is given by $\text{TP},\text{FP}],[\text{FN},\text{TN}]]=[[66,45],[2,119]]$.
c. Human response time assessment
Figure 10: Response time determination in the VR experiment
Figure 11: Survey results from the VR experiment
Would issuing warnings to pedestrians help reduce users’ response time and increase their safety awareness? To test this hypothesis, we designed virtual reality (VR) experiments in Unity3D [144], where warnings are provided via voice and text displayed on the VR headset Meta Quest 3. There are two traffic scenarios, one involving an interaction between the participant (i.e., a pedestrian) and an oncoming scooter, and the other between the participant and an oncoming vehicle. In each scenario, participants first receive no warning and then a warning, and their response times are recorded. As illustrated in Fig. 10, when traffic approaches at time $t$, the user either receives a warning or does not. The user stops after a response time $t_{r}$, which is then recorded. The response time distributions are presented in Fig. 12, with the ‘no warning’ condition shown in yellow and the ‘warning’ condition in blue. The issuance of warning has reduced people’s average response time by $0.62\mathrm{s}$ (in Scenario 1) and $1.11\mathrm{s}$ (in Scenario 2), respectively. The demographics of the participants are shown in the upper section of Fig. 11. The bar chart summarizes the participants’ survey responses regarding VR and traffic realism, the perceived importance of safety warnings, and comfort within the VR environment, with mean and standard deviation indicated on the top left corners of each subfigure.
Figure 12: Comparison of response times with and without a warning. A scooter (in Scenario 1) and a vehicle (in Scenario 2) poses the danger, respectively.
IV-C3 Discussions
We would like to highlight that our NYC’s testbed deployment demonstrates generalizability. As the biggest metropolis of the U.S., NYC’s dense population, limited space, and multimodal transportation systems (consisting of public transit with buses and subway, driving, ferry) motivate us to study the interactions between automobiles and VRUs (including pedestrians, cyclists, and surging micromoblity and e-mobility). Located in one of the world’s busiest regions - Manhattan of the NYC, the COSMOS testbed offers a natural laboratory for us to collect abundant data and conduct live experiments, which would otherwise not be possible.
We leverage the COSMOS testbed to design and implement sensor (e.g., camera, LiDAR) data analytics systems and evaluate them in urban settings to ensure they can scale to thousands of connected intersections. To achieve this, we ensure that the designed analytics systems are efficient in resource consumption, particularly network bandwidth and GPU usage. We also ensure that these systems are modular and deployable in a distributed manner across multiple edge servers and cloud infrastructure. Fig. 3 demonstrates such a modular and distributable system. This design enables flexibility, allowing the system to be deployed in settings where a single edge server cannot handle the full workload, requiring distribution across several edge servers. Therefore, the system must be designed such that each module can be easily deployed on a different server and can seamlessly communicate with other modules, regardless of whether they reside on the same server. Additionally, the AI models used must be replaceable or generalizable to accommodate different environments, camera angles, heights, and other variations. With these principles in mind, we use the COSMOS testbed to develop scalable systems.
The NYC’s dense high-rise buildings enable us to leverage legacy infrastructure for traffic sensing and monitoring, and facilitates the instrumentation of cameras and communication nodes. Note that there is a growing trend of studies relying on emerging communication protocols like C-V2X for vehicle DTs [145, 146], which requires customized infrastructure and user-end devices. For rapid and scalable deployment, our pilot experiments utilize contemporary protocols, which hold the potential to further leverage the city’s legacy traffic cameras and wireless communication infrastructure for scalability and cost-effectiveness. Our technologies could be replaced by emerging methods that can achieve lower latency and higher bandwidths. For instance, software defined features of the COSMOS testbed and the flexibility of radios make it possible to deploy emerging technologies such as C-V2X and experiment with novel low-latency applications. Since major intersections in the neighborhood has similar sensor suites, it would not need additional infrastructure deployment to generalize our system to network wide multi-intersection settings.
We will extend our use cases to multi-pedestrian warning scenarios. Smartphone‑to‑cloud beaconing with an Adaptive Multi‑Mode (AMM) scheme enables real‑time location sharing from multiple devices, allowing the server to issue timely risk alerts [147]. A Vehicle‑to‑Pedestrian (V2P) system [148] leveraging Bluetooth Low Energy (BLE) advertising on smartphones broadcasts standardized Personal Safety Messages (PSM) to address the multi‑device communication challenge in road safety. Moreover, efficient multi-device communication is essential for real-time data sharing and coordination among mobile users in smart urban environments. Key challenges include ensuring reliable connectivity in dense settings, minimizing communication latency, and maintaining user privacy. Advanced techniques such as device-to-device (D2D) communication [149], opportunistic networking [150], and privacy-preserving data aggregation [151] have been developed to address these issues.
V Conclusions and open questions
In this paper, we first review the AI methods applied to every stage of the DT pipeline, from object detection, tracking, prediction, simulation, to traffic operation and management. Then, leveraging the unique characteristics of NYC, we propose a DT architecture and present a safety-critical use case, namely, intersection safety warning. Three evaluation methods are performed, including measuring latency, simulating collision reduction, and VR experiments for human responses.
A growing number of sensors, explosive amounts of data, and increasing computational powers have opened up tremendous opportunities for researchers to apply AI to create, train, evaluate, and streamline DT pipelines for uban traffic management. Subsequently, we will present emerging trends, challenges, and open questions for the development of DT.
V-A Emerging trends